# Pretrained Reversible Generation as Unsupervised Visual Representation Learning

### v1

@@ B.1 Training Process Details
Tabs. B1 and B2 list the hyperparameters used for both pretraining and finetuning across the CIFAR-10, Tiny-ImageNet, and ImageNet datasets.
@@
Table B1: Experimental settings across datasets for pretraining.
@@
Table B2: Experimental settings across datasets for finetuning.
To enhance the reproducibility of results across various multi-stage and multi-GPU experiments, we calculate the learning rate using Eq. B1.
LR $=\text{LR}_{\text{base}}\times\frac{\text{num processes}\times\text{Batch Sizee}}{512}$ (B1)
@@ B.2 Evaluation of Training Efficiency
Most research experiments, including the main experiments and ablation studies, are completed within 2 hours to 3 days. To demonstrate the training efficiency, Tab. B3 report the per-epoch runtime for each dataset used in our experiments.
@@
Table B3: Time per epoch during the training process.* means that the model is trained on $4*8$ A100 GPUs.$t_{\text{span}}$ represents the total sampling length, while $t_{\text{cutoff}}$ indicates the point along the trajectory where fine-tuning begins. For example, $20/10$ means a trajectory spanning $20$ steps from $x_{0}$ to $x_{1}$, with fine-tuning starting from the midpoint of the trajectory.
B.3 Details of Out-of-Distribution Experiments
@@
PGD [madry2017towards] $51.08$ $33.46$ $\downarrow 17.62$ $24.00$ $\downarrow 27.08$
PLAT [kireev2022effectiveness] $51.29$ $37.92$ $\downarrow 13.37$ $29.05$ $\downarrow 22.24$
@@
RSE [liu2018towards] $53.74$ $27.99$ $\downarrow 25.75$ $18.92$ $\downarrow 34.82$
ENResNet [wang2020enresnet] $49.26$ $25.83$ $\downarrow 23.43$ $19.01$ $\downarrow 30.25$
@@
AugMix [hendrycks2019augmix] $52.82$ $37.74$ $\downarrow 15.08$ $28.66$ $\downarrow 24.16$
AutoAug [cubuk2019autoaugment] $52.63$ $35.14$ $\downarrow 17.49$ $25.36$ $\downarrow 27.27$
@@
Table B4: (OOD: extrapolation datasets) Performance on Tiny-ImageNet-C.Averge represents the accuracy across all corruption levels, with corruption severity ranging from $1$ to $5$.
There is no direct correspondence between the test images of Tiny ImageNet and Tiny ImageNet-C, and the images in Tiny ImageNet-C do not overlap with the training images of Tiny ImageNet. We report the comparative results on Tiny ImageNet-C in Tab. B4.
Figure B1: Reverse Generation Process on the Swiss Roll Dataset.
Figure B2: Reverse Generation Process on the TinyImageNet val set.
B.4 Reverse Generation Process
Fig. B1 illustrates the reverse generation process from $x_{0}$ to $x_{1}$ after finetuning. Furthermore, Figs. B2 and B3 present the reverse generation results on the TinyImageNet dataset after pretraining and finetuning, respectively. Finally, Fig. B4 demonstrates the reverse process before and after applying fog corruption to the images.
Appendix C Ablation Studies
C.1 Loss Type
Tab. C1 shows that label smoothing reduces overconfident predictions, improves model calibration, and enhances robustness.
Dataset LabelSmooth Loss Cross-Entropy Loss
CIFAR-$10$ $97.59$ $96.18$
TinyImageNet $71.12$ $70.15$
Table C1: Comparison of LabelSmooth Loss and Cross-Entropy Loss on different datasets.
C.2 The Number of Timeteps
Our findings in Tab. C2 show that longer time spans generally lead to better accuracy. On CIFAR-10 with the ICFM flow model, a $t$-span of 10 achieves accuracy comparable to the best result at $t=100$. In contrast, TinyImageNet requires a $t$-span of 15 to achieve similar performance.
@@ C.2 The Number of Timeteps
Table C2: Comparison of Model Performance Over Varying Time Spans.
Figure B3: Reverse Generation Process on the TinyImageNet train set.The first row represents the fully pretrained reverse generative process, the second row shows the reverse generative process after extensive finetuning.
Figure B4: Reverse Generation Process on the TinyImageNet-C Dataset. The first row represents the fully pretrained reverse generative process, the second row shows the reverse process after extensive finetuning, and the third row illustrates the reverse generative process under fog corruption.

### v2

@@ B.1 Training Process Details
Tabs. 1 and 2 list the hyperparameters used for both pre-training and fine-tuning across the CIFAR-10, Tiny-ImageNet, and ImageNet datasets.
@@
Table 1: Experimental settings across datasets for pre-training.
@@
Table 2: Experimental settings across datasets for fine-tuning.
To enhance the reproducibility of results across various multi-stage and multi-GPU experiments, we calculate the learning rate using Eq. 1.
LR $=\text{LR}_{\text{base}}\times\frac{\text{num processes}\times\text{Batch Size}}{512}$ (1)
@@ B.2 Evaluation of Training Efficiency
Most research experiments, including the main experiments and ablation studies, are completed within several hours. To demonstrate training efficiency concretely, Tab. 3 reports the run-time per epoch for each data set used in our experiments.
@@
Table 3: Mean time cost per epoch during the training process. * means that the model is trained on $4*8$ A100 GPUs. $t_{\text{span}}$ represents the total sampling length, while $t_{\text{cutoff}}$ indicates the point along the trajectory where fine-tuning begins. For example, $20/10$ means a trajectory spanning $20$ steps from $x_{0}$ to $x_{1}$, with fine-tuning starting from the midpoint of the trajectory.
B.3 Evaluation of Inference Efficiency
Table 4 presents the inference efficiency of our method on the CIFAR-10 and Tiny-ImageNet test sets.
Model CIFAR-10 (s) Tiny-ImageNet (s)
PRG-GVP-S 4 10
PRG-ICFM-S 3 9
PRG-OTCFM-S 3 8
Table 4: Mean inference time per epoch on the CIFAR-10 and Tiny-ImageNet test datasets.
Appendix C Ablation Studies
C.1 Loss Type
Tab. 5 shows the results of different loss types. Compared to the standard cross entropy loss, label smoothing reduces overconfident predictions, improves model calibration, and improves robustness.
Dataset LabelSmooth Loss Cross-Entropy Loss
CIFAR-$10$ $97.59$ $96.18$
TinyImageNet $71.12$ $70.15$
Table 5: (Loss Type) Comparison of LabelSmooth Loss and Cross-Entropy Loss on different datasets.
C.2 ODE Solver Type
During fine-tuning, we evaluated different ODE solvers for the reverse process: Euler (first-order), Midpoint (second-order via midpoint evaluations), RK4 (fourth-order Runge-Kutta), and Dopri5 (adaptive step sizes with a fifth-order method). Tabs. 6 and 7 compares their performance on the Cifar-10, Tiny-ImageNet dataset. The results show no significant performance differences, underscoring the method’s consistent effectiveness across various solvers.
solver_type Euler Midpoint RK4 Dopri5
OTCFM $97.65$ $97.63$ $97.66$ $97.72$
ICFM $97.59$ $97.55$ $97.56$ $97.61$
GVP $97.35$ $97.30$ $97.32$ $97.41$
Table 6: (ODE Solver) Performance of various solvers on Cifar-10. Different solvers don’t yield obvious differences.
Solver Type Euler Midpoint RK4 Dopri5
PRG-OTCFM $71.33$ $71.29$ $71.30$ $71.36$
PRG-ICFM $71.12$ $71.11$ $71.13$ $71.23$
PRG-GVP $70.89$ $70.99$ $70.84$ $70.85$
Table 7: (ODE Solver) Performance of various solvers on Tiny-ImageNet. Different solvers don’t yield obvious differences.
C.3 Details of Out-of-Distribution Experiments
@@
PGD [43] $51.08$ $33.46$ $\downarrow 17.62$ $24.00$ $\downarrow 27.08$
PLAT [31] $51.29$ $37.92$ $\downarrow 13.37$ $29.05$ $\downarrow 22.24$
@@
RSE [37] $53.74$ $27.99$ $\downarrow 25.75$ $18.92$ $\downarrow 34.82$
ENResNet [64] $49.26$ $25.83$ $\downarrow 23.43$ $19.01$ $\downarrow 30.25$
@@
AugMix [25] $52.82$ $37.74$ $\downarrow 15.08$ $28.66$ $\downarrow 24.16$
AutoAug [15] $52.63$ $35.14$ $\downarrow 17.49$ $25.36$ $\downarrow 27.27$
@@
Table 8: (OOD: extrapolated datasets) Performance on Tiny-ImageNet-C.Averge represents the accuracy across all corruption levels, with corruption severity ranging from $1$ to $5$.
There is no direct correspondence between the test images of Tiny ImageNet and Tiny ImageNet-C, and the images in Tiny ImageNet-C do not overlap with the training images of Tiny ImageNet. We report the comparative results on Tiny ImageNet-C in Tab. 8.
C.4 The Number of Timesteps
Our findings in Tab. 9 show that longer time spans generally lead to better accuracy. On CIFAR-10 with the ICFM flow model, a $t$-span of 10 achieves accuracy comparable to the best result at $t=100$. In contrast, TinyImageNet requires a $t$-span of 15 to achieve similar performance.
@@ C.4 The Number of Timesteps
Table 9: Comparison of Performance Over Varying Time Spans.
C.5 Transfer Experiment on Dual-Task Datasets
Following the approach described in Sec. 4.4.3, we transfer the powerful text-to-image model SiT-XL [42] from the community. However, instead of directly applying it to downstream classification tasks, we leverage its diffusion trajectory for feature extraction across multiple tasks. Specifically, we set the diffusion trajectory to 30 steps, extracting features from the 10-th reverse diffusion step for CIFAR classification and the 20-th step for Tiny-ImageNet.
Our algorithm, PRG-SiT-XL2, achieves an accuracy of 96.9% on CIFAR and 70.2% on Tiny-ImageNet.
Figure 1: Reverse Generation Process on the Swiss Roll Dataset. Each color represents a different class. After diffusion, samples from the same class become more clustered, while the previously unoccupied white space, corresponding to out-of-class regions, is pushed outward.
Figure 2: Reverse Generation Process on the TinyImageNet val set. The first row represents the fully pretrained reverse generative process, the second row shows the reverse generative process after extensive fine-tuning.
Appendix D Reverse Generation Process
Fig. 1 illustrates the reverse generation process from $x_{0}$ to $x_{1}$ after fine-tuning. Furthermore, Figs. 2 and 3 present the reverse generation results on the TinyImageNet dataset after pre-training and fine-tuning, respectively. Finally, Fig. 4 demonstrates the reverse process before and after applying fog corruption to the images.
Figure 3: Reverse Generation Process on the TinyImageNet train set. The first row represents the fully pretrained reverse generative process, the second row shows the reverse generative process after extensive fine-tuning.
Figure 4: Reverse Generation Process on the TinyImageNet-C Dataset. The first row represents the fully pretrained reverse generative process, the second row shows the reverse process after extensive fine-tuning, and the third row illustrates the reverse generative process under fog corruption.